# Mutual Information Surprise: Rethinking Unexpectedness in Autonomous Systems

### v1

@@ Abstract
Recent breakthroughs in autonomous experimentation have demonstrated remarkable physical capabilities, yet their cognitive control remains limited—often relying on static heuristics or classical optimization. A core limitation is the absence of a principled mechanism to detect and adapt to the unexpectedness. While traditional surprise measures—such as Shannon or Bayesian Surprise—offer momentary detection of deviation, they fail to capture whether a system is truly learning and adapting. In this work, we introduce Mutual Information Surprise (MIS), a new framework that redefines surprise not as anomaly detection, but as a signal of epistemic growth. MIS quantifies the impact of new observations on mutual information, enabling autonomous systems to reflect on their learning progression. We develop a statistical test sequence to detect meaningful shifts in estimated mutual information and propose a mutual information surprise reaction policy (MISRP) that dynamically governs system behavior through sampling adjustment and process forking. Empirical evaluations—on both synthetic domains and a dynamic pollution map estimation task—show that MISRP-governed strategies significantly outperform classical surprise-based approaches in stability, responsiveness, and predictive accuracy. By shifting surprise from reactive to reflective, MIS offers a path toward more self-aware and adaptive autonomous systems.
@@ 1 Introduction
This shift in perspective aligns with formal definitions of surprise in information theory and computational psychology, such as Shannon surprise  (?), Bayesian surprise  (?), Bayes Factor surprise  (?), and Confidence-Corrected surprise  (?). These surprise definitions quantify unexpectedness by modeling deviations from prior beliefs or probability distributions. In the following section, we will delve deeper into these existing measures and evaluate whether they truly serve the intended role of identifying opportunities, as human surprise does, more than merely flagging anomalies. Using current surprise definitions,  (?) demonstrated that treating surprising events not as noise to be removed but as catalysts for learning can significantly enhance a system’s learning speed. Additional empirical evidence shows that incorporating surprise as a learning mechanism can improve autonomy in domains such as autonomous driving  (?, ?, ?) and manufacturing  (?, ?).
In our research, we find that existing definitions of surprise require significant improvement. Their close resemblance to anomaly detection measures suggests that they may not effectively support higher levels of autonomy. Specifically, a robust surprise measure should emphasize knowledge acquisition and adaptability, rather than treating unexpectedness merely as a deviation from the norm—an approach that current surprise definitions tend to adopt. We therefore argue that it is essential to develop a novel surprise metric that inherently fosters learning and deepens an autonomous system’s understanding of the underlying processes it encounters. To capture this dynamic capability, we introduce the Mutual Information Surprise (MIS)—a new framework that redefines how autonomous systems interpret and respond to unexpected events. MIS quantifies the degree of both frustration and enlightenment associated with new observations, measuring their impact on refining the system’s internal understanding of its environment. We also demonstrate the differences that arise when applying mutual information surprise, as opposed to relying solely on classical surprise definitions, highlighting MIS’s potential to meaningfully enhance autonomous learning and decision-making.
@@ 2 Current Surprise Definitions and Their Limitations
Classical definitions of surprise, such as Shannon and Bayesian Surprise, provide elegant mathematical frameworks for quantifying unexpectedness. However, these approaches often fall short in capturing the core mechanisms driving adaptive behavior: continuous learning and flexible model updating. This section revisits and analyzes existing formulations, elaborating on their conceptual foundations and outlining both their strengths and limitations.
@@
The family of Shannon Surprise metrics emphasizes the improbability of observed data, typically independent of explicit model parameters. This class broadly aligns with “observation” and “probabilistic-mismatch” surprises as categorized in  (?). The central question which the Shannon family of surprises tries to answer is straightforward: How unlikely is the observation?
The most widely recognized measure is Shannon Surprise  (?), formally defined as:
@@
interpreting surprise directly through event rarity. Although conceptually clear and mathematically elegant, this definition has a significant limitation: encountering a Shannon Surprise does not inherently imply knowledge acquisition. Consider, for instance, a uniform dartboard—a stochastic yet entirely understood system. Each outcome has an equally low probability, thus appearing “surprising” under Shannon’s definition, despite humans neither genuinely finding these outcomes surprising nor gaining any additional knowledge by observing them. In other words, the focus of Shannon Surprise is statistical rarity rather than genuine knowledge gain.
@@
$S_{\text{Residual}}(\mathbf{x})=\large|\underset{\mathbf{x}^{\prime}}{\min}\{-\log p(\mathbf{x}^{\prime})\}-(-\log p(\mathbf{x}))\large|=\underset{\mathbf{x}^{\prime}}{\max}\log p(\mathbf{x}^{\prime})-\log p(\mathbf{x}).$
In the dartboard example, Residual Information Surprise becomes zero for all outcomes, as $p(\mathbf{x}^{\prime})$ remains constant for every $\mathbf{x}^{\prime}$, accurately reflecting an absence of genuine surprise. However, this formulation introduces a conceptual challenge, as determining $\underset{\mathbf{x}^{\prime}}{\max}\log p(\mathbf{x}^{\prime})$ implicitly presumes an omniscient oracle, an assumption typically infeasible in practice.
@@
$S_{\text{Shannon}}(\mathbf{x})\propto\|\mathbf{x}-\mu_{\mathbf{x}}\|_{2}^{2},$
thus linking surprise with deviation from the mean. Similarly, assuming a Laplace distribution recovers an absolute error interpretation, termed Absolute Error Surprise in  (?):
$S_{\text{Shannon}}(\mathbf{x})\propto|\mathbf{x}-\mu_{\mathbf{x}}|.$
We note that both Squared Error Surprise and Absolute Error Surprise are commonly utilized metrics in anomaly detection literature  (?, ?, ?).
@@
Bayesian Surprises, by contrast, explicitly model belief updates. These measures quantify the degree to which a new observation alters the internal model, shifting the focus from event rarity to epistemic impact. This concept parallels the “belief-mismatch” surprise in the taxonomy by  (?).
The canonical formulation, introduced in  (?), defines Bayesian Surprise as the Kullback–Leibler divergence between the prior and posterior distributions over parameters:
$S_{\text{Bayes}}(\mathbf{x})=D_{\text{KL}}\left(\pi(\boldsymbol{\theta}\mid\mathbf{x})\,\|\,\pi(\boldsymbol{\theta})\right).$
This measure offers a principled approach to belief revision and naturally aligns with learning mechanisms. In theory, it encourages agents to reduce surprise through model updates, providing a pathway toward adaptive autonomy.
@@
$S_{\text{CC}}(\mathbf{x})=D_{\text{KL}}\left(\pi(\boldsymbol{\theta})\,\|\,\pi^{\prime}(\boldsymbol{\theta}\mid\mathbf{x})\right),$
where $\pi^{\prime}(\boldsymbol{\theta}\mid\mathbf{x})$ represents the updated belief assuming a uniform prior. This confidence-corrected formulation remains sensitive to new data irrespective of prior history. In the $J$ particle example, employing Confidence-Corrected Surprise would trigger a genuine surprise, as the posterior remains responsive to the novel observation without the inertia introduced by extensive historical data.
@@
$S_{\text{BF}}(\mathbf{x})=\frac{p(\mathbf{x}\mid\pi^{0}(\boldsymbol{\theta}))}{p(\mathbf{x}\mid\pi^{t}(\boldsymbol{\theta}))},$
where $\pi^{0}(\boldsymbol{\theta})$ represents the naïve (untrained) prior and $\pi^{t}(\boldsymbol{\theta})$ the informed belief based on all prior observations up to time $t$ (before observing $\mathbf{x}$). This ratio quantifies how strongly the current observation supports the naïve prior over the informed prior. In practice, the effectiveness of both Confidence-Corrected and Bayes Factor Surprises heavily depends on constructing appropriate priors—a task often challenging and subjective.
@@
$S_{\text{Postdictive}}(\mathbf{x})=D_{\text{KL}}\left(P(\mathbf{y}\mid\boldsymbol{\theta}^{\prime},\mathbf{x})\,\|\,P(\mathbf{y}\mid\boldsymbol{\theta},\mathbf{x})\right),$ (2)
where $\boldsymbol{\theta}$ and $\boldsymbol{\theta}^{\prime}$ denote parameters before and after the update, respectively.  (?) argue that computing KL divergence in the output space is more computationally tractable for variational models but potentially less expressive when output variance depends on the input (e.g., under heteroskedastic conditions).
@@
These categorizations are not strictly delineated. For instance, Residual Information Surprise incorporates a conceptual element common to the Bayesian Surprise family—providing a baseline against which the observed data is contrasted with. On the other hand, Bayes Factor Surprise, despite being explicitly Bayesian in its formulation, closely resembles a Shannon Surprise conditioned on alternative priors. Furthermore, notwithstanding their philosophical distinctions, Bayesian and Shannon Surprises often behave similarly in practice; we provide further details on this observation in Section 4.
@@
Both surprise families share other critical limitations: they are single-instance measures and inherently one-sided measures. Being single instance means that they assess surprise based solely on the marginal impact of individual observations, without explicitly modeling cumulative learning dynamics over time, whereas being one-sided means that they have a decision threshold on one single side, offering limited expressiveness since human perceptions of surprise can range from positive to negative.
@@ 3 Mutual Information Surprise
In this section, we introduce the concept of Mutual Information Surprise (MIS). We first explore the intuition and motivation underlying this concept, followed by the development of a novel, theoretically grounded testing sequence. We then discuss the implications when this test sequence is violated and propose a reaction policy contingent on different types of violations. Table 1 summarizes the perspective differences between Mutual Information Surprise vs Shannon and Bayesian family of surprises.
Shannon Family
Bayesian Family
MIS
3.1 What Do We Expect from a Surprise?
In human cognition, surprise often triggers reflection and adaptation. A computational analog should similarly prompt deeper examination and enhanced understanding, transcending mere statistical rarity and indicating an opportunity for learning.
To formalize this perspective, consider a system governed by a functional mapping $f:\mathbf{x}\rightarrow\mathbf{y}$, with observations drawn from a joint distribution $P(\mathbf{x},\mathbf{y})$. This system is well-regulated, meaning the input distribution $P(\mathbf{x})$, output distribution $P(\mathbf{y})$, and joint distribution $P(\mathbf{x},\mathbf{y})$ are time-invariant. This definition expands the traditional notion of time-invariance by explicitly including consistent exposure $P(\mathbf{x})$, aligning closely with human trust in persistent patterns across rules and experiences.
To quantify system understanding, we use mutual information (MI)  (?), defined as
$I(\mathbf{x},\mathbf{y})=\mathbb{E}_{\mathbf{x},\mathbf{y}}\left[\log\frac{p(\mathbf{y}\mid\mathbf{x})}{p(\mathbf{y})}\right]=H(\mathbf{x})+H(\mathbf{y})-H(\mathbf{x},\mathbf{y})=H(\mathbf{y})-H(\mathbf{y}\mid\mathbf{x}),$ (3)
where $H(\cdot)$ denotes entropy, measuring uncertainty or chaos of a random variable. Mutual information quantifies the reduction in uncertainty about $\mathbf{y}$ given knowledge of $\mathbf{x}$. A high $I(\mathbf{x},\mathbf{y})$ indicates strong comprehension of $f$, whereas stagnation or a decrease in $I(\mathbf{x},\mathbf{y})$ signals stalled learning. For the aforementioned well-regulated system, $I(\mathbf{x},\mathbf{y})$ remains constant.
Typically, mutual information $I(\mathbf{x},\mathbf{y})$ is estimated via maximum likelihood estimation (MLE)  (?); details of the MLE estimator are provided in the Appendix. Empirical estimation of $I(\mathbf{x},\mathbf{y})$ is, however, downward biased for clean data with a low noise level  (?):
$\mathbb{E}[\hat{I}(\mathbf{x},\mathbf{y})]\leq I(\mathbf{x},\mathbf{y}).$
Interestingly, this bias can serve as an informative feature: As experience accumulates, $\mathbb{E}[\hat{I}(\mathbf{x},\mathbf{y})]$ should increase and approach the true value $I(\mathbf{x},\mathbf{y})$, determined by $p(\mathbf{x})$ and function $f$. Thus, a monotonic growth in mutual information estimate signals learning.
Returning to our core question—what do we expect from a surprise? Unlike classical surprise measures (Shannon or Bayesian), which focus narrowly on conditional distributions and rarity, we posit that a surprise measure should reflect whether learning occurred. Noticing the connection between mutual information growth and learning, we define surprise as a deviation from expected mutual information growth. Specifically, we define Mutual Information Surprise (MIS) as the difference in mutual information estimates after incorporating new observations:

### v3

@@ Abstract
A community of researchers appears to think that a machine can be surprised and have introduced various surprise measures, principally the Shannon Surprise and the Bayesian Surprise. The questions of what constitutes a surprise and how to react to one still elicit debates. In this work, we introduce Mutual Information Surprise (MIS), a new framework that redefines surprise not as anomaly measure, but as a signal of epistemic growth. Furthermore, we develop a statistical test sequence that could trigger a surprise reaction and propose a MIS-based reaction policy that dynamically governs system behavior through sampling adjustment and process forking. Empirical evaluations—on both synthetic domains and a dynamic pollution map estimation task—show that a system governed by the MIS-based reaction policy significantly outperforms those under classical surprise-based approaches in stability, responsiveness, and predictive accuracy. The important implication of our new proposal is that MIS quantifies the impact of new observations on mutual information, shifts surprise from reactive to reflective, enables reflection on learning progression, and thus offers a path toward self-aware and adaptive autonomous systems. We expect the new surprise measure to play a critical role in further advancing autonomous systems on their ability to learn and adapt in a complex and dynamic environment.
@@ 1 Introduction
This shift in perspective aligns with formal definitions of surprise in information theory and computational psychology, such as Shannon surprise (?), Bayesian surprise (?), Bayes Factor surprise (?), and Confidence-Corrected surprise (?). These surprise definitions quantify unexpectedness by modeling deviations from prior beliefs or probability distributions. Using some of the existing surprise definitions, (?) demonstrated that treating surprising events not as noise to be removed but as catalysts for learning can significantly enhance a system’s learning speed. Additional empirical evidence shows that incorporating surprise as a learning mechanism can improve autonomy in domains such as autonomous driving (?, ?, ?) and manufacturing (?, ?). In Section 2, we will delve deeper into these existing measures and evaluate whether they truly serve the intended role of identifying opportunities, as human surprise does, more than merely flagging anomalies.
In this paper, we consider a general class of input-output systems described by a functional mapping, $f(\cdot)$, such that
$f:\mathbf{x}\in\mathcal{X}\rightarrow\mathbf{y}\in\mathcal{Y},$ (1)
where $\mathbf{x}$ denotes the inputs and $\mathbf{y}$ denotes the outputs to the system. The data pair, $\{\mathbf{x},\mathbf{y}\}$, is drawn independently (not necessarily identical) from a potentially non-stationary joint distribution $P(\mathbf{x},\mathbf{y})$.
Most existing anomaly and surprise measures implicitly assume that the underlying distribution $P(\mathbf{x},\mathbf{y})$ remains stationary over time. Such an assumption is restrictive: truly autonomous systems operate in evolving environments, where system dynamics, external conditions, and even objectives may shift. A surprise measure should not merely detect deviations under a fixed model, but rather actively capture potential changes in both inputs and the systems.
Motivated by the need to address this limitation, we argue that it is essential to develop a new surprise measure that does not anchor itself to the stationarity of $P(\mathbf{x},\mathbf{y})$, but instead inherently fosters learning and deepens an autonomous system’s understanding of the underlying processes it encounters. To capture this dynamic capability, we introduce the Mutual Information Surprise (MIS)—a new framework that redefines how autonomous systems interpret and respond to unexpected events. MIS quantifies the degree of both frustration and enlightenment associated with new observations, measuring their impact on refining the system’s internal understanding of its environment. We contrast the outcomes made by applying mutual information surprise with that by relying solely on classical surprise definitions to highlight MIS’s potential to enhance learning and decision-making.
@@ 2 Current Surprise Definitions and Their Limitations
Classical definitions of surprise, such as Shannon and Bayesian Surprise, provide elegant mathematical frameworks for quantifying unexpectedness. Indeed, the notion of surprise plays a central role in the field of active inference (?), where it serves as a key driver of perception and action. However, these approaches often fall short of capturing the mechanisms that underlie adaptive behavior, namely continuous learning and flexible model updating. In this section, we revisit existing formulations of surprise, examining their conceptual foundations while highlighting both their strengths and their limitations.
@@
The family of Shannon Surprise metrics emphasizes the improbability of observed data, typically independent of explicit model parameters. This class broadly aligns with “observation” and “probabilistic-mismatch” surprises as categorized in (?). The central question which the Shannon family of surprises tries to answer is: How unlikely is the observation? For that a Shannon Surprise (?) is commonly defined as:
@@
interpreting surprise directly through event rarity. The above expression defines Shannon surprise in terms of the marginal input observation $\mathbf{x}$. It is not difficult to see that the same definition is applicable to outputs $\mathbf{y}$. For a practical system with an input-output structure as expressed in Equation (1), surprise is more naturally defined with respect to the predictive distribution of outputs given inputs, namely in a conditional perspective. This is to say, instead of evaluating $\log p(\mathbf{x})$, one considers
$S_{\text{Shannon}}(\mathbf{x},\mathbf{y})=-\log p(\mathbf{y}\mid\mathbf{x})$ (3)
for Shannon Surprise. In this work, we adopt this conditional viewpoint when evaluating the Shannon family of surprise measures.
Although conceptually clear and mathematically elegant, the above definition has a significant limitation: encountering a Shannon Surprise does not inherently imply knowledge acquisition. Consider, for instance, a uniform dartboard—a stochastic yet entirely understood system. Each outcome has an equally low probability, thus appearing “surprising” under Shannon’s definition, despite humans neither genuinely finding these outcomes surprising nor gaining any additional knowledge by observing them. In other words, the focus of Shannon Surprise is statistical rarity rather than genuine knowledge gain.
@@
$S_{\text{Residual}}(\mathbf{x},\mathbf{y})=\large|\underset{\mathbf{y}^{\prime}}{\min}\{-\log p(\mathbf{y}^{\prime}|\mathbf{x})\}-(-\log p(\mathbf{y}|\mathbf{x}))\large|=\underset{\mathbf{y}^{\prime}}{\max}\log p(\mathbf{y}^{\prime}|\mathbf{x})-\log p(\mathbf{y}|\mathbf{x}).$
In the dartboard example, Residual Information Surprise becomes zero for all outcomes, as $p(\mathbf{y}^{\prime}|\mathbf{x})$ remains constant for every outcome $\mathbf{y}^{\prime}$, accurately reflecting an absence of genuine surprise. However, this formulation introduces a conceptual challenge, as determining $\underset{\mathbf{y}^{\prime}}{\max}\log p(\mathbf{y}^{\prime}|\mathbf{x})$ implicitly presumes an omniscient oracle, an assumption typically infeasible in practice.
@@
$S_{\text{Shannon}}(\mathbf{x},\mathbf{y})\propto\|\mathbf{y}-\boldsymbol{\mu}_{\mathbf{y}\mid\mathbf{x}}\|_{2}^{2},$
thus linking surprise with deviation from the mean. Similarly, assuming a Laplace distribution, Shannon surprise recovers an absolute error interpretation, termed Absolute Error Surprise in (?):
$S_{\text{Shannon}}(\mathbf{x},\mathbf{y})\propto|\mathbf{y}-\boldsymbol{\mu}_{\mathbf{y}\mid\mathbf{x}}|.$
Both Squared Error Surprise and Absolute Error Surprise are commonly utilized metrics in anomaly detection literature (?, ?, ?).
@@
Bayesian Surprises, by contrast, explicitly model belief updates. Rather than assessing the improbability of an observation, the Bayesian family assumes that the data-generating process is governed by an underlying parameter $\boldsymbol{\theta}$, and defines surprise in terms of how a new observation changes the posterior belief over $\boldsymbol{\theta}$. These measures quantify the degree to which a new observation alters the internal model, shifting the focus from event rarity to epistemic impact. This concept parallels the “belief-mismatch” surprise in the taxonomy by (?).
The canonical formulation, introduced in (?), defines Bayesian Surprise as the Kullback-Leibler divergence between the prior and posterior distributions over parameters:
$S_{\text{Bayes}}(\mathbf{x})=D_{\text{KL}}\left(\pi(\boldsymbol{\theta}\mid\mathbf{x})\,\|\,\pi(\boldsymbol{\theta})\right).$ (4)
The above definition is presented in terms of $\mathbf{x}$. Like in the case of Shannon surprise, the Bayesian family of surprise measures naturally extend to input-output systems as well. This extension follows from recognizing that the parameter $\boldsymbol{\theta}$ characterizes the functional relationship between the input $\mathbf{x}$ and the output $\mathbf{y}$. As such, observations arrive as paired data $(\mathbf{x},\mathbf{y})$, and belief updates occur over the parameters governing the conditional mapping $P(\mathbf{y}\mid\mathbf{x},\boldsymbol{\theta})$. Consequently, the Bayesian Surprise for input-output systems can be written as
$S_{\text{Bayes}}(\mathbf{x},\mathbf{y})=D_{\text{KL}}\left(\pi(\boldsymbol{\theta}\mid\mathbf{x},\mathbf{y})\,\|\,\pi(\boldsymbol{\theta})\right),$ (5)
where the posterior $\pi(\boldsymbol{\theta}\mid\mathbf{x},\mathbf{y})$ reflects the updated belief after observing the input-output pair.
The Bayesian Surprise measure offers a principled approach to belief revision and naturally aligns with learning mechanisms. In theory, it encourages agents to reduce surprise through model updates, providing a pathway toward adaptive autonomy.
@@
$S_{\text{CC}}(\mathbf{x},\mathbf{y})=D_{\text{KL}}\left(\pi(\boldsymbol{\theta})\,\|\,\pi^{\prime}(\boldsymbol{\theta}\mid\mathbf{x},\mathbf{y})\right),$
where $\pi^{\prime}(\boldsymbol{\theta}\mid\mathbf{x},\mathbf{y})$ represents the updated belief assuming a uniform prior. This confidence-corrected formulation remains sensitive to new data irrespective of prior history. In the $J$ particle example, employing Confidence-Corrected Surprise would trigger a genuine surprise, as the posterior remains responsive to the novel observation without the inertia introduced by extensive historical data.
@@
$S_{\text{BF}}(\mathbf{x},\mathbf{y})=\frac{p(\mathbf{x},\mathbf{y}\mid\pi^{0}(\boldsymbol{\theta}))}{p(\mathbf{x},\mathbf{y}\mid\pi^{t}(\boldsymbol{\theta}))},$
where $\pi^{0}(\boldsymbol{\theta})$ represents the naïve (untrained) prior and $\pi^{t}(\boldsymbol{\theta})$ the informed belief based on all prior observations up to time $t$ (before observing $(\mathbf{x},\mathbf{y})$). This ratio quantifies how strongly the current observation supports the naïve prior over the informed prior. In practice, the effectiveness of both Confidence-Corrected and Bayes Factor Surprises heavily depends on constructing appropriate priors—a task often challenging and subjective.
@@
$S_{\text{Postdictive}}(\mathbf{x},\mathbf{y})=D_{\text{KL}}\left(P(\mathbf{y}\mid\boldsymbol{\theta}^{\prime},\mathbf{x})\,\|\,P(\mathbf{y}\mid\boldsymbol{\theta},\mathbf{x})\right),$ (6)
where $\boldsymbol{\theta}$ and $\boldsymbol{\theta}^{\prime}$ denote the predictive model parameters before and after the update, respectively. (?) argue that computing KL divergence in the output space is more computationally tractable for variational models but potentially less expressive when output variance depends on the input (e.g., under heteroskedastic conditions).
@@
These categorizations are not strictly delineated. For instance, Residual Information Surprise incorporates a conceptual element common to the Bayesian Surprise family—providing a baseline against which the observed data is contrasted with. On the other hand, Bayes Factor Surprise, despite being explicitly Bayesian in its formulation, closely resembles a Shannon Surprise conditioned on alternative priors. Notwithstanding their philosophical distinctions, Bayesian and Shannon Surprises often behave similarly in practice; we provide further details on this observation in Section 4.
@@
Both surprise families share other critical limitations: they are single-instance measures by design and are inherently one-sided measures. Being single instance means that they primarily focus on surprise based solely on the marginal impact of individual observations, without explicitly modeling cumulative learning dynamics over time, whereas being one-sided means that they have a decision threshold on one single side, offering limited expressiveness since human perceptions of surprise can range from positive to negative.
@@ 3 Mutual Information Surprise
In this section, we introduce the concept of Mutual Information Surprise (MIS). We first explore the intuition and motivation underlying this concept, followed by the development of a novel, theoretically grounded testing sequence. We then discuss the implications when this test sequence is violated and propose a reaction policy contingent on different types of violations.
3.1 An Alternative Surprise Definition
Recall that we consider input-output systems described by a functional mapping $f(\cdot):\mathbf{x}\rightarrow\mathbf{y}$, with observations drawn from a joint distribution $P(\mathbf{x},\mathbf{y})$. Classical surprise measures are implicitly constructed under the assumption that $P(\mathbf{x},\mathbf{y})$ remains stationary over time. However, autonomous engineering systems often operate in evolving environments where such stationarity assumptions may not hold. What this means is that not all changes in $P(\mathbf{x},\mathbf{y})$ are of interest or should trigger a surprise reaction. This prompts us to look for a new tool to define surprise for the input-output system we are dealing with.
Our research leads us to the concept of mutual information (MI), which was introduced by (?) (although the specific term, “mutual information,” was coined by Robert Fano nearly ten year later) and defined as
$I(\mathbf{x};\mathbf{y})=\mathbb{E}_{\mathbf{x},\mathbf{y}}\left[\log\frac{p(\mathbf{y}\mid\mathbf{x})}{p(\mathbf{y})}\right]=H(\mathbf{x})+H(\mathbf{y})-H(\mathbf{x},\mathbf{y})=H(\mathbf{y})-H(\mathbf{y}\mid\mathbf{x}),$ (7)
where $H(\cdot)$ denotes entropy, which measures the uncertainty of a random variable. The mutual information, $I(\mathbf{x};\mathbf{y})$, which is always non-negative, quantifies the reduction in uncertainty about $\mathbf{y}$ when $\mathbf{x}$ is observed. In fact, mutual information can serve as a quantitative tool to benchmark “system learning” because Fano’s inequality (?) establishes that independently of prediction models, the best possible learning performance is governed by mutual information. Specifically, an increasing $I(\mathbf{x};\mathbf{y})$ indicates that the system is gaining in understanding of the underlying functional relationship $f(\cdot)$, whereas stagnation or decline in $I(\mathbf{x};\mathbf{y})$ suggests that the learning has stalled. Importantly, not all changes in $P(\mathbf{x},\mathbf{y})$ will cause a change in mutual information, a property that we look for to accommodate dynamic autonomous systems.
Unlike the traditional change detection literature which usually defines the null hypothesis (the baseline) as a stationary in-control distribution, we define our null hypothesis as an input-output system with a stationary mutual information. We formalize this null hypothesis through the following notion of a well-regulated autonomous system:
Definition 1.
A well-regulated input-output autonomous system may exhibit non-stationary joint distributions $P(\mathbf{x},\mathbf{y})$ over time, yet its mutual information $I(\mathbf{x};\mathbf{y})$ remains unchanged.
For an autonomous system to trigger a surprise, it is when the mutual information of the system changes, meaning that the system is out of the well-regulated region it has been operating under. In practice, the mutual information $I(\mathbf{x};\mathbf{y})$ is typically estimated using maximum likelihood estimation (MLE) (?); some details regarding the estimator used in this work are provided in the Appendix. The estimation introduces uncertainty, so that statistical decision thresholds need to be established as a function of the prescribed significant level; we will provide the upper/lower statistical decision thresholds in Section 3.2.
The above arguments lead us to introduce a Mutual Information Surprise (MIS), defined as the change in estimated mutual information after incorporating new observations: